Could please help me with this question?
The function that relates the Cost ($) of an oil well to its Depth (x) in feet is as follows:
C(x) = -146,983 + 58.72 * x
Find and interpret C(3500).

Answer by ankor@dixie-net.com(12700)

The function that relates the Cost ($) of an oil well to its Depth (x) in feet is as follows:
C(x) = -146,983 + 58.72 * x
Find and interpret C(3500).
:
Substitute 3500 for x.
C(3500) = -146983 + 58.72(3500)
C(3500) = -146983 + 205520
C(3500) = $58,537 cost to drill to 3500 ft